OpenStudy (anonymous):

A man kicks a soccer ball with a speed of 14 m/s at an angle of 40 degree (A)- how long does it take for the b all to reach the soccer goal if the young man is 9.85m away from it ?

OpenStudy (anonymous):

I forgot what the equation is...help

OpenStudy (anonymous):



OpenStudy (anonymous):

x= 9.85 = 14cos40 *t

OpenStudy (anonymous):

so x is t which is 9.85?

OpenStudy (anonymous):

the equation I'm using or was is V^2sin*2(40)/g

OpenStudy (anonymous):

I'm still lost on it

OpenStudy (anonymous):

I just need an equation and might of put the wrong one

OpenStudy (anonymous):

x=V*t

OpenStudy (anonymous):

\[x=x _{0}+V _{0x}t+\frac{ 1 }{ 2 }a _{x} t ^{2}\] from the previous image, we see the origin, therefore xo is 0, and there is no acceleration in the x direction, therefore 1/2 at^2 is zero so x = Vcost theta * t

OpenStudy (anonymous):

yes, yu can also use t= V^2Sin theta over g it shud give the same answer

OpenStudy (anonymous):

okay well I just want to see if its right

OpenStudy (anonymous):

yes it is,

OpenStudy (anonymous):

I'm gonna solve it and give you the answer just see if I'm right or wrong

OpenStudy (anonymous):

I got 14.9...yes or no?

OpenStudy (anonymous):

I did 14^2 which is 196sin (40) / 9.8

OpenStudy (anonymous):

so 9.8 seconds?

OpenStudy (anonymous):

no, it is 0.92 seconds, use the \[t = \frac{ 9.85 }{ 14\cos 40 }\]

OpenStudy (anonymous):

ahhh! I did it backward ugh

OpenStudy (anonymous):

I don't see how I'm getting -.469

OpenStudy (anonymous):

in my calculator i put in 9.85 / 14cos(40)

OpenStudy (anonymous):

make sure your calculator is in degree mode,

OpenStudy (anonymous):

https://www.desmos.com/calculator, yu can use that, make sure the settings is in degree mode

OpenStudy (anonymous):

yeah well I'm putting even in putting in degree mode I got .539

OpenStudy (anonymous):

I'm gonna try it on an online calculator

OpenStudy (anonymous):

good well I got it on here well I need help solving another